∀n ≥ 1, Xn i=1 i(i!) = (n + 1)! − 1


1
Expert's answer
2021-11-08T16:02:20-0500

"\displaystyle{\sum^n_{i=1} }i(i!)=(n+1)!-1"


proof by induction:


for n=1:

"\displaystyle{\sum^1_{i=1} }1\cdot(1!)=1"

"(1+1)!-1=1"


let for n=k:


"\displaystyle{\sum^k_{i=1} }i(i!)=(k+1)!-1"


then for n=k+1:


"\displaystyle{\sum^{k+1}_{i=1} }i(i!)=\displaystyle{\sum^{k}_{i=1} }i(i!)+(k+1)(k+1)!=(k+1)!-1+(k+1)(k+1)!="


"=(k+2)(k+1)!-1=(k+2)!-1"